# LLM Unlearning with LLM Beliefs

### v1

@@ Appendix E Further Experimental Setup
This section consolidates the experimental setup and protocols. Appendix E.1 introduces the datasets/benchmarks and any preprocessing or split strategy. Appendix E.2 specifies the evaluation metrics and aggregation protocol used across datasets. Appendix E.3 details implementation and training settings. Appendix E.4 lists hyperparameter settings and search spaces for all baselines and our methods.
@@ E.1 Benchmarks
TOFU (Task of Fictitious Unlearning). TOFU (Maini et al., 2024) introduces a controlled setup for studying LLM unlearning by constructing a synthetic question-answering dataset about fictitious authors. Each author profile includes 20 Q&A pairs generated via GPT-4, incorporating diverse personal details such as birthplace, genre, and literary awards. Crucially, as the data is hallucinated, it guarantees no prior exposure during pretraining, offering a clean testbed to isolate the effects of fine-tuning and unlearning. The dataset is divided into forget and retain splits (e.g., 10%, 5%, or 1% forget), enabling granular control over unlearning difficulty. TOFU is particularly useful for probing semantic-level forgetting in scenarios where the model’s sole exposure to a concept is via fine-tuning, and any residual generation after unlearning signals incomplete forgetting. This makes it ideal for detecting spurious retention due to memorization or semantic generalization.
@@
WMDP (Weapons of Mass Destruction Proxy). WMDP (Li et al., 2024b) is motivated by dual-use risk mitigation in LLMs. It comprises 3,668 expert-written multiple-choice questions spanning biosecurity, cybersecurity, and chemical safety. Each question tests whether the model retains potentially dangerous knowledge that could be exploited for malicious purposes, such as synthesizing toxins or executing cyberattacks. Rather than aiming to remove specific documents or identities, WMDP focuses on domain-level unlearning, where the goal is to suppress model capabilities related to hazardous information while preserving broader knowledge in adjacent fields (e.g., general biology or programming). The benchmark is particularly relevant for closed-source deployment, where models must remain secure even when subject to jailbreak or adversarial finetuning. WMDP provides a public, open-access proxy for red-teaming evaluations previously restricted to private labs, thus enabling community-wide research on hazard-aware unlearning.
@@ E.3 Implementation Setup
Training Configuration. Our experiments mainly follow the configurations in OpenUnlearning (Dorna et al., 2025) and Wang et al. (2025d); Yang et al. (2025). Specifically, we employ the AdamW optimizer (Loshchilov & Hutter, 2019) across all benchmarks. For TOFU, we set the batch size to 32, learning rate to 1e-5, training for 10 epochs with a linear scheduler, 1 epoch of warm-up, and weight decay of 0.01. For MUSE, we use a constant learning rate scheduler with a batch size of 32, a learning rate of 1e-5, and train for 10 epochs. Results are selected from 10 checkpoints saved at each epoch, following Yang et al. (2025). For WMDP, we adopt a batch size of 16, learning rate of 4e-6, and train for 125 steps with 25 steps for warm-up, following Yang et al. (2025).

### v2

@@ D.1 Learning Dynamics
D.3 Proof of Theorem 5.3
See 5.3
Proof.
We work at a single step $t\to t+1$ of SGD on $\mathcal{L}_{\rm BSS}^{\textnormal{off}}({\bm{\theta}};\bm{\mathsf{\chi}}_{\rm u})$ with learning rate $\eta$. By definition of the off-policy BS-S objective,
${\bm{\theta}}_{t+1}={\bm{\theta}}_{t}-\eta\nabla_{\bm{\theta}}\mathcal{L}_{\rm BSS}^{\textnormal{off}}({\bm{\theta}}_{t};\bm{\mathsf{\chi}}_{\rm u})={\bm{\theta}}_{t}-\eta\sum_{m=0}^{N}\omega_{m}\nabla_{\bm{\theta}}\mathcal{L}_{\rm BST}({\bm{\theta}}_{t};\bm{\mathsf{\chi}}_{\rm u}^{m}),$ (11)
where the augmented pairs $\{\bm{\mathsf{\chi}}_{\rm u}^{m}\}_{m=0}^{N}$ and weights $\{\omega_{m}\}_{m=0}^{N}$ are fixed under the off-policy construction.
Fix any observing pair $\bm{\mathsf{\chi}}_{\rm o}=[{\mathbf{x}}_{\rm u};{\mathbf{y}}_{\rm o}]$. Linearizing the logit network ${\mathbf{z}}(\bm{\mathsf{\chi}}_{\rm o};{\bm{\theta}})$ around ${\bm{\theta}}_{t}$ in the lazy eNTK regime (as in Lem. 5.1) gives
$\Delta{\mathbf{z}}(\bm{\mathsf{\chi}}_{\rm o})\;\equiv\;{\mathbf{z}}(\bm{\mathsf{\chi}}_{\rm o};{\bm{\theta}}_{t+1})-{\mathbf{z}}(\bm{\mathsf{\chi}}_{\rm o};{\bm{\theta}}_{t})=\nabla_{\bm{\theta}}{\mathbf{z}}(\bm{\mathsf{\chi}}_{\rm o};{\bm{\theta}}_{t})\,({\bm{\theta}}_{t+1}-{\bm{\theta}}_{t})+\mathcal{O}(\eta^{2}).$ (12)
Substituting Eq. (11) into Eq. (12) and using linearity of the gradient,
$\Delta{\mathbf{z}}(\bm{\mathsf{\chi}}_{\rm o})$ $=-\,\eta\,\nabla_{\bm{\theta}}{\mathbf{z}}(\bm{\mathsf{\chi}}_{\rm o};{\bm{\theta}}_{t})\sum_{m=0}^{N}\omega_{m}\nabla_{\bm{\theta}}\mathcal{L}_{\rm BST}({\bm{\theta}}_{t};\bm{\mathsf{\chi}}_{\rm u}^{m})+\mathcal{O}(\eta^{2})$
$=-\,\eta\sum_{m=0}^{N}\omega_{m}\nabla_{\bm{\theta}}{\mathbf{z}}(\bm{\mathsf{\chi}}_{\rm o};{\bm{\theta}}_{t})\,\nabla_{\bm{\theta}}\mathcal{L}_{\rm BST}({\bm{\theta}}_{t};\bm{\mathsf{\chi}}_{\rm u}^{m})+\mathcal{O}(\eta^{2}).$ (13)
For each $m$, the BS-T loss depends on $\bm{\theta}$} only through the logits ${\mathbf{z}}(\bm{\mathsf{\chi}}_{\rm u}^{m};{\bm{\theta}})$, so by the chain rule and the definition of the BS-T residual,
$\nabla_{\bm{\theta}}\mathcal{L}_{\rm BST}({\bm{\theta}}_{t};\bm{\mathsf{\chi}}_{\rm u}^{m})=\nabla_{\bm{\theta}}^{\top}{\mathbf{z}}(\bm{\mathsf{\chi}}_{\rm u}^{m};{\bm{\theta}}_{t})\,\mathcal{G}_{{\rm BST},t}(\bm{\mathsf{\chi}}_{\rm u}^{m}),$
where $\mathcal{G}_{{\rm BST},t}(\bm{\mathsf{\chi}}_{\rm u}^{m})=\nabla_{\mathbf{z}}\mathcal{L}_{\rm BST}(\bm{\mathsf{\chi}}_{\rm u}^{m})\big|_{{\bm{\theta}}_{t}}$ and its token-wise components coincide with $\mathcal{G}^{i}_{\rm BST}$ in Thm. 5.2. Plugging this into Eq. (13) yields
$\Delta{\mathbf{z}}(\bm{\mathsf{\chi}}_{\rm o})$ $=-\,\eta\sum_{m=0}^{N}\omega_{m}\nabla_{\bm{\theta}}{\mathbf{z}}(\bm{\mathsf{\chi}}_{\rm o};{\bm{\theta}}_{t})\,\nabla_{\bm{\theta}}^{\top}{\mathbf{z}}(\bm{\mathsf{\chi}}_{\rm u}^{m};{\bm{\theta}}_{t})\,\mathcal{G}_{{\rm BST},t}(\bm{\mathsf{\chi}}_{\rm u}^{m})+\mathcal{O}(\eta^{2})$
$=-\,\eta\sum_{m=0}^{N}\omega_{m}\,\mathcal{K}_{t}(\bm{\mathsf{\chi}}_{\rm o},\bm{\mathsf{\chi}}_{\rm u}^{m})\,\mathcal{G}_{{\rm BST},t}(\bm{\mathsf{\chi}}_{\rm u}^{m})+\mathcal{O}(\eta^{2}),$ (14)
where $\mathcal{K}_{t}(\bm{\mathsf{\chi}}_{\rm o},\bm{\mathsf{\chi}}_{\rm u}^{m})=\nabla_{\bm{\theta}}{\mathbf{z}}(\bm{\mathsf{\chi}}_{\rm o};{\bm{\theta}}_{t})\nabla_{\bm{\theta}}^{\top}{\mathbf{z}}(\bm{\mathsf{\chi}}_{\rm u}^{m};{\bm{\theta}}_{t})$ is the empirical eNTK.
The Jacobian of $\log\pi_{{\bm{\theta}}_{t}}(\cdot\mid\bm{\mathsf{\chi}}_{\rm o})$ with respect to ${\mathbf{z}}(\bm{\mathsf{\chi}}_{\rm o})$ is exactly the same as in Lem. 5.1, namely
$\mathcal{A}_{t}(\bm{\mathsf{\chi}}_{\rm o})\;\equiv\;\nabla_{\mathbf{z}}\log\pi_{{\bm{\theta}}_{t}}(\cdot\mid\bm{\mathsf{\chi}}_{\rm o})=\bm{I}-\mathbbm{1}\,\pi_{{\bm{\theta}}_{t}}^{\top}(\cdot\mid\bm{\mathsf{\chi}}_{\rm o}).$
Applying the chain rule to Eq. (14), we obtain
$\Delta\log\pi_{t}(\cdot\mid\bm{\mathsf{\chi}}_{\rm o})=\mathcal{A}_{t}(\bm{\mathsf{\chi}}_{\rm o})\,\Delta{\mathbf{z}}(\bm{\mathsf{\chi}}_{\rm o})=-\,\eta\,\mathcal{A}_{t}(\bm{\mathsf{\chi}}_{\rm o})\sum_{m=0}^{N}\omega_{m}\mathcal{K}_{t}(\bm{\mathsf{\chi}}_{\rm o},\bm{\mathsf{\chi}}_{\rm u}^{m})\,\mathcal{G}_{{\rm BST},t}(\bm{\mathsf{\chi}}_{\rm u}^{m})+\mathcal{O}(\eta^{2}).$
Taking the component corresponding to the candidate response ${\mathbf{y}}_{\rm o}$ yields the claimed update
$\Delta\log\pi_{t}({\mathbf{y}}_{\rm o}\mid\bm{\mathsf{\chi}}_{\rm o})=-\,\eta\,\mathcal{A}_{t}(\bm{\mathsf{\chi}}_{\rm o})\sum_{m=0}^{N}\omega_{m}\mathcal{K}_{t}(\bm{\mathsf{\chi}}_{\rm o},\bm{\mathsf{\chi}}_{\rm u}^{m})\,\mathcal{G}_{{\rm BST},t}(\bm{\mathsf{\chi}}_{\rm u}^{m})+\mathcal{O}(\eta^{2}),$
which matches the statement of Thm. 5.3. ∎
D.4 Discussion on On-Policy BS-S
For completeness, we now clarify why the learning-dynamics analysis in Thm. 5.3 does not directly extend to the on-policy BS-S variant used in our implementation.
On-policy BS-S objective.
Recall that off-policy BS-S fixes an augmented set of sequences $\{\bm{\mathsf{\chi}}_{\rm u}^{m}\}_{m=0}^{N}$ in advance and optimizes
$\mathcal{L}_{\rm BSS}^{\textnormal{off}}({\bm{\theta}};{\mathbf{x}}_{\rm u})=\sum_{m=0}^{N}\omega_{m}\,\mathcal{L}_{\rm BST}({\bm{\theta}};\bm{\mathsf{\chi}}_{\rm u}^{m}),$ (15)
which fits the fixed-data assumption in Lem. 5.1. In contrast, the on-policy BS-S objective for a single forget prompt ${\mathbf{x}}_{\rm u}$ can be written as
$\mathcal{L}_{\rm BSS}^{\textnormal{on}}({\bm{\theta}};{\mathbf{x}}_{\rm u})=(1-\lambda_{\rm BSS})\,\mathcal{L}_{\rm BST}({\bm{\theta}};\bm{\mathsf{\chi}}_{\rm u}^{0})+\lambda_{\rm BSS}\,\mathbb{E}_{\tilde{{\mathbf{y}}}_{\rm u}\sim\mu_{{\bm{\theta}}}(\cdot\mid{\mathbf{x}}_{\rm u})}\bigl[\mathcal{L}_{\rm BST}({\bm{\theta}};[{\mathbf{x}}_{\rm u};\tilde{{\mathbf{y}}}_{\rm u}])\bigr],$ (16)
where $\bm{\mathsf{\chi}}_{\rm u}^{0}=[{\mathbf{x}}_{\rm u};{\mathbf{y}}_{\rm u}]$ is the original forget pair and $\mu_{{\bm{\theta}}}(\cdot\mid{\mathbf{x}}_{\rm u})$ denotes the on-policy sampling distribution induced by the current model $\pi_{{\bm{\theta}}}$ on the forget prompt ${\mathbf{x}}_{\rm u}$ (e.g., given by autoregressive decoding with temperature and nucleus sampling). Conditioned on a particular sampled sequence $\tilde{{\mathbf{y}}}_{\rm u}$, the inner $\mathcal{L}_{\rm BST}({\bm{\theta}};[{\mathbf{x}}_{\rm u};\tilde{{\mathbf{y}}}_{\rm u}])$ is still evaluated under teacher forcing on the fixed sequence ${\mathbf{x}}_{\rm u};\tilde{{\mathbf{y}}}_{\rm u}]$.
Gradient decomposition.
Differentiating Eq. (16) with respect to $\bm{\theta}$} gives
$\nabla_{{\bm{\theta}}}\mathcal{L}_{\rm BSS}^{\textnormal{on}}({\bm{\theta}};{\mathbf{x}}_{\rm u})$ $=(1-\lambda_{\rm BSS})\,\nabla_{{\bm{\theta}}}\mathcal{L}_{\rm BST}({\bm{\theta}};\bm{\mathsf{\chi}}_{\rm u}^{0})+\lambda_{\rm BSS}\,\nabla_{{\bm{\theta}}}\mathbb{E}_{\tilde{{\mathbf{y}}}_{\rm u}\sim\mu_{{\bm{\theta}}}(\cdot\mid{\mathbf{x}}_{\rm u})}\bigl[\mathcal{L}_{\rm BST}({\bm{\theta}};[{\mathbf{x}}_{\rm u};\tilde{{\mathbf{y}}}_{\rm u}])\bigr].$ (17)
For the expectation term, we use the standard score-function (log-derivative) identity. Let
$f({\bm{\theta}},\tilde{{\mathbf{y}}}_{\rm u})\;=\;\mathcal{L}_{\rm BST}({\bm{\theta}};[{\mathbf{x}}_{\rm u};\tilde{{\mathbf{y}}}_{\rm u}]).$
Then
$\nabla_{{\bm{\theta}}}\mathbb{E}_{\tilde{{\mathbf{y}}}_{\rm u}\sim\mu_{{\bm{\theta}}}}\bigl[f({\bm{\theta}},\tilde{{\mathbf{y}}}_{\rm u})\bigr]$ $=\mathbb{E}_{\tilde{{\mathbf{y}}}_{\rm u}\sim\mu_{{\bm{\theta}}}}\bigl[\nabla_{{\bm{\theta}}}f({\bm{\theta}},\tilde{{\mathbf{y}}}_{\rm u})\bigr]+\mathbb{E}_{\tilde{{\mathbf{y}}}_{\rm u}\sim\mu_{{\bm{\theta}}}}\bigl[f({\bm{\theta}},\tilde{{\mathbf{y}}}_{\rm u})\,\nabla_{{\bm{\theta}}}\log\mu_{{\bm{\theta}}}(\tilde{{\mathbf{y}}}_{\rm u}\mid{\mathbf{x}}_{\rm u})\bigr].$ (18)
Substituting Eq. (18) back into Eq. (17) yields
$\nabla_{{\bm{\theta}}}\mathcal{L}_{\rm BSS}^{\textnormal{on}}({\bm{\theta}};{\mathbf{x}}_{\rm u})$ $=(1-\lambda_{\rm BSS})\,\nabla_{{\bm{\theta}}}\mathcal{L}_{\rm BST}({\bm{\theta}};\bm{\mathsf{\chi}}_{\rm u}^{0})$
$\quad+\lambda_{\rm BSS}\,\mathbb{E}_{\tilde{{\mathbf{y}}}_{\rm u}\sim\mu_{{\bm{\theta}}}}\bigl[\nabla_{{\bm{\theta}}}\mathcal{L}_{\rm BST}({\bm{\theta}};[{\mathbf{x}}_{\rm u};\tilde{{\mathbf{y}}}_{\rm u}])\bigr]$ (19)
$\quad+\lambda_{\rm BSS}\,\mathbb{E}_{\tilde{{\mathbf{y}}}_{\rm u}\sim\mu_{{\bm{\theta}}}}\bigl[\mathcal{L}_{\rm BST}({\bm{\theta}};[{\mathbf{x}}_{\rm u};\tilde{{\mathbf{y}}}_{\rm u}])\,\nabla_{{\bm{\theta}}}\log\mu_{{\bm{\theta}}}(\tilde{{\mathbf{y}}}_{\rm u}\mid{\mathbf{x}}_{\rm u})\bigr].$
The first two terms in RHS of Eq. (19) have the same form as in the off-policy case: they are expectations of BS-T gradients on teacher-forced pairs ${\mathbf{x}}_{\rm u};{\mathbf{y}}_{\rm u}]$ and ${\mathbf{x}}_{\rm u};\tilde{{\mathbf{y}}}_{\rm u}]$. Conditioned on $\tilde{{\mathbf{y}}}_{\rm u}$, each term can be written as
$\nabla_{{\bm{\theta}}}\mathcal{L}_{\rm BST}({\bm{\theta}};\bm{\mathsf{\chi}})=\nabla_{{\bm{\theta}}}^{\top}{\mathbf{z}}(\bm{\mathsf{\chi}};{\bm{\theta}})\,\mathcal{G}_{{\rm BST}}(\bm{\mathsf{\chi}}),$
so they admit an AKG-style kernel decomposition exactly as in Appx. D.1D.3.
The third term in RHS of Eq. (19) is qualitatively different: it is a policy-gradient-like term that explicitly involves $\nabla_{{\bm{\theta}}}\log\mu_{{\bm{\theta}}}(\tilde{{\mathbf{y}}}_{\rm u}\mid{\mathbf{x}}_{\rm u})$. Expanding the latter under the autoregressive factorization,
$\mu_{{\bm{\theta}}}(\tilde{{\mathbf{y}}}_{\rm u}\mid{\mathbf{x}}_{\rm u})=\prod_{i=1}^{|\tilde{{\mathbf{y}}}_{\rm u}|}\pi_{{\bm{\theta}}}(\tilde{{\mathbf{y}}}_{\rm u}^{i}\mid{\mathbf{x}}_{\rm u},\tilde{{\mathbf{y}}}_{\rm u}^{<i}),$
gives
$\nabla_{{\bm{\theta}}}\log\mu_{{\bm{\theta}}}(\tilde{{\mathbf{y}}}_{\rm u}\mid{\mathbf{x}}_{\rm u})=\sum_{i=1}^{|\tilde{{\mathbf{y}}}_{\rm u}|}\nabla_{{\bm{\theta}}}\log\pi_{{\bm{\theta}}}(\tilde{{\mathbf{y}}}_{\rm u}^{i}\mid{\mathbf{x}}_{\rm u},\tilde{{\mathbf{y}}}_{\rm u}^{<i}),$ (20)
where each conditional distribution is evaluated on the sampled prefix $\tilde{{\mathbf{y}}}_{\rm u}^{<i}$ produced by the current model rather than on a dataset prefix under teacher forcing.
Why the AKG framework does not apply.
The AKG decomposition in Lem. 5.1 (and the subsequent Thms. 5.2 and 5.3) relies on two structural conditions:
  1. 1.
    The training examples (pairs $\bm{\mathsf{\chi}}$) are fixed and independent of $\bm{\theta}$} during the SGD step; all parameter dependence enters only through the logit network ${\mathbf{z}}(\bm{\mathsf{\chi}};{\bm{\theta}})$ evaluated under teacher forcing on these fixed sequences.
  2. 2.
    The total gradient can therefore be written as a sum of terms of the form $\nabla_{{\bm{\theta}}}^{\top}{\mathbf{z}}(\bm{\mathsf{\chi}};{\bm{\theta}})\mathcal{G}(\bm{\mathsf{\chi}})$, for some residuals $\mathcal{G}(\bm{\mathsf{\chi}})$ defined on this fixed dataset.
For the on-policy BS-S objective Eq. (16), the first two lines of Eq. (19) satisfy these conditions conditionally on the sampled sequences $\tilde{{\mathbf{y}}}_{\rm u}$: they are averages of BS-T gradients on teacher-forced pairs ${\mathbf{x}}_{\rm u};{\mathbf{y}}_{\rm u}]$ and ${\mathbf{x}}_{\rm u};\tilde{{\mathbf{y}}}_{\rm u}]$. If we were to ignore the dependence of $\mu_{{\bm{\theta}}}$ on $\bm{\theta}$} (i.e., drop the third line), the resulting approximation would reduce exactly to the off-policy analysis in Thm. 5.3.
However, the policy-gradient term in the third line of Eq. (19) cannot be written as a finite sum of the form
$\sum_{m}\tilde{\omega}_{m}\,\nabla_{{\bm{\theta}}}^{\top}{\mathbf{z}}(\bm{\mathsf{\chi}}^{m};{\bm{\theta}})\tilde{\mathcal{G}}(\bm{\mathsf{\chi}}^{m})$
over a fixed collection of teacher-forced sequences $\{\bm{\mathsf{\chi}}^{m}\}$. Indeed, Eq. (20) involves gradients of log-probabilities evaluated along on-policy trajectories whose prefixes $\tilde{{\mathbf{y}}}_{\rm u}^{<i}$ are themselves functions of $\bm{\theta}$}. As $\bm{\theta}$} changes, both the support and the relative weights of $\mu_{{\bm{\theta}}}(\cdot\mid{\mathbf{x}}_{\rm u})$ change, so there is no parameter-independent dataset that captures all the sequences contributing to the score term.
Consequently, the overall update on the logits ${\mathbf{z}}(\bm{\mathsf{\chi}}_{\rm o};{\bm{\theta}})$ induced by $\mathcal{L}_{\rm BSS}^{\textnormal{on}}$ contains an additional component that cannot be expressed through the kernel $\mathcal{K}_{t}(\cdot,\cdot)$ and a residual evaluated on a fixed training set, and the AKG-based learning-dynamics derivation in Appx. D.1D.3 does not apply. From a modeling perspective, this is analogous to the distinction between supervised finetuning on a fixed corpus and on-policy RLHF methods: the latter require a separate analysis that jointly tracks both the parameter updates and the evolving sampling distribution $\mu_{{\bm{\theta}}}$. Developing such an RL-style learning-dynamics theory for on-policy BS-S is beyond the scope of this work and we leave it for future research.
@@ Appendix E Further Experimental Setup
This section consolidates the experimental setup and protocols. §E.1 introduces the datasets/benchmarks and any preprocessing or split strategy. §E.2 specifies the evaluation metrics and aggregation protocol used across datasets. §E.3 details implementation and training settings. §E.4 lists hyperparameter settings and search spaces for all baselines and our methods.
@@ E.1 Benchmarks
TOFU (Task of Fictitious Unlearning). TOFU (Maini et al., 2024) introduces a controlled setup for studying LLM unlearning by constructing a synthetic question-answering dataset about fictitious authors. Each author profile includes 20 Q&A pairs generated via GPT-4, incorporating diverse personal details such as birthplace, genre, and literary awards. Crucially, as the data is hallucinated, it guarantees no prior exposure during pre-training, offering a clean testbed to isolate the effects of fine-tuning and unlearning. The dataset is divided into forget and retain splits (e.g., 10%, 5%, or 1% forget), enabling granular control over unlearning difficulty. TOFU is particularly useful for probing semantic-level forgetting in scenarios where the model’s sole exposure to a concept is via fine-tuning, and any residual generation after unlearning signals incomplete forgetting. This makes it ideal for detecting spurious retention due to memorization or semantic generalization.
@@
WMDP (Weapons of Mass Destruction Proxy). WMDP (Li et al., 2024b) is motivated by dual-use risk mitigation in LLMs. It comprises 3,668 expert-written multiple-choice questions spanning biosecurity, cybersecurity, and chemical safety. Each question tests whether the model retains potentially dangerous knowledge that could be exploited for malicious purposes, such as synthesizing toxins or executing cyberattacks. Rather than aiming to remove specific documents or identities, WMDP focuses on domain-level unlearning Huang et al. (2023; 2024; 2025), where the goal is to suppress model capabilities related to hazardous information while preserving broader knowledge in adjacent fields (e.g., general biology or programming). The benchmark is particularly relevant for closed-source deployment, where models must remain secure even when subject to jailbreak or adversarial finetuning. WMDP provides a public, open-access proxy for red-teaming evaluations previously restricted to private labs, thus enabling community-wide research on hazard-aware unlearning.
@@ E.3 Implementation Setup
Training Configuration. Our experiments mainly follow the configurations in OpenUnlearning (Dorna et al., 2025) and Wang et al. (2025d); Yang et al. (2025). Specifically, we employ the AdamW optimizer (Loshchilov and Hutter, 2019) across all benchmarks. For TOFU, we set the batch size to 32, learning rate to 1e-5, training for 10 epochs with a linear scheduler, 1 epoch of warm-up, and weight decay of 0.01. For MUSE, we use a constant learning rate scheduler with a batch size of 32, a learning rate of 1e-5, and train for 10 epochs. Results are selected from 10 checkpoints saved at each epoch, following Yang et al. (2025). For WMDP, we adopt a batch size of 16, learning rate of 4e-6, and train for 125 steps with 25 steps for warm-up, following Yang et al. (2025).